Given a hypothetical consumption function of the form:
Y = C + I0 + G0
C = α + β
Where: = Y – T
Y = Income
T = Taxes
Compute the equilibrium level of income and consumption
1
Expert's answer
2018-07-09T15:25:08-0400
If Y = C + I0 + G0, C = α + β(Y - T),
then the equilibrium level of income will be Y = α + β(Y - T) = I0 + G0 and consumption will be C = α + β(Y - T) = Y - I0 - G0